What is the mmimum uncertainty in the position of a neutron if the uncertainty in its speed is 0.0250 m/s?
Answer #1

According to the Uncertainty principle

Δx * Δp = h/4* π (h= planck's constant = 6.626 *10-34J.s)

Neutron mass (m) = 1.675 * 10-27 Kg

Given Uncertainty in speed (Δv) = 0.0250 m/s

Δp = Δv * m = 0.0250 m/s * 1.675 * 10-27 Kg = 4.1875 * 10-29 Kg.m/s

Δx * Δp = h/4* π

Δx = 6.626 *10-34J.s /(4*(22/7)*Δp) (1J = Kg.m2/s2)

Δx = 6.626 *10-34J.s /(4*(22/7)* 4.1875 * 10-29 Kg.m/s ) = 1.2586 * 10-6 m

The minimum Uncertainty in the position of the neutron is equal to 1.2586 * 10-6 m